OpenStudy (anonymous):

Solve the system of equations by graphing. 7x + 7y = 0 5x − 2y = 14 (x, y) = ( )

OpenStudy (anonymous):

Change the equations into slope intercept form: y = mx + b That will make them easy to draw on a graph.

OpenStudy (anonymous):

how do I do that?

OpenStudy (anonymous):

Move everything but the 'y' to the right side of the equation: 7x + 7y = 0 7y = -7x now divide both sides by 7: y = -x This is in the form: y = mx + b with m = -1 and b = 0 b is the y - intercept (where it crosses the y axis, so it is where x = 0) So the line goes through the point (0,0) with a slope, the value for m, of -1 A slope of -1 means from (0,0), the next point on the line will be down 1 and to the right 1, or at the point (1,-1). Do the same thing with the other equation. Then you can see on your graph where the two lines intersect. That will the x and y values that solve both of the equations.

OpenStudy (anonymous):

When you do this to the second equation, the value for m (the slope) will be 5/2. From the point that you find for b, which is the value for y where it crosses the y axis, your next point on the line, with a slope of 5/2, will be up 5 and to the right 2.